OpenStudy (anonymous):

The measure of a single angle in a regular octagon is (4x + 12)°. Find the value of x. Show equations and all work that leads to your answer.

jimthompson5910 (jim_thompson5910):

If A is the interior angle of a regular polygon, then A = 180(n-2)/n Plug in n = 8 to get A = 180(8-2)/8 A = 135 So an octagon has each interior angle of 135 degrees

jimthompson5910 (jim_thompson5910):

each interior angle is also equal to 4x+12 degrees, so... 4x + 12 = 135 now solve for x

OpenStudy (anonymous):

x would be 8.43?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

I am so confused lol

jimthompson5910 (jim_thompson5910):

4x + 12 = 135 4x = 135-12 4x = 123 x = ??

OpenStudy (anonymous):

30.75?

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

x = 123/4 or x = 30.75

OpenStudy (anonymous):

thanks so much!

jimthompson5910 (jim_thompson5910):

yw